# Addressing the Infinite Variance Problem in Fermionic Monte Carlo Simulations: Retrospective Error Remediation and the Exact Bridge Link Method

### v1

@@ III.3 Application to the SU(4) Hubbard model: charge-$4e$ correlations
We next present a real application of the exact bridge link method to a challenging problem, namely the attractive SU(4) Hubbard model. This model is of particular interest for studying quartet condensation, also known as charge-$4e$ superconductivity [52, 53, 54, 55]. Unlike the conventional charge-$2e$ superconducting state, characterized by a pair order parameter involving two creation operators, the charge-$4e$ state is described by the quartet order parameter. Computing the charge-$4e$ correlation function is challenging due to both infinite variance and a large noise-to-signal ratio — particularly because the long-range correlation signal may remain small even in the ordered phase. As we illustrate below, the new method makes it possible to perform “clean” simulations to tackle this problem.
@@ V Acknowledgments
$\begin{array}[]{llll}\beta=\frac{c^{+}-c^{-}}{c^{+}+c^{-}}&,a_{n}=\left(\frac{2\Gamma(\alpha)\sin(\frac{\pi\alpha}{2})}{\pi(c^{+}+c^{-})}\right)^{1/\alpha}n^{-1/\alpha}&,b_{n}=0&\text{if \ \ }0<\alpha<1,\\ \beta=\frac{c^{+}-c^{-}}{c^{+}+c^{-}}&,a_{n}=\left(\frac{2\Gamma(\alpha)\sin(\frac{\pi\alpha}{2})}{\pi(c^{+}+c^{-})}\right)^{1/\alpha}n^{-1/\alpha}&,b_{n}=n\mathrm{Im}\log\Phi_{X}(a_{n})&\text{if \ \ }\alpha=1,\\ \beta=\frac{c^{+}-c^{-}}{c^{+}+c^{-}}&,a_{n}=\left(\frac{2\Gamma(\alpha)\sin(\frac{\pi\alpha}{2})}{\pi(c^{+}+c^{-})}\right)^{1/\alpha}n^{-1/\alpha}&,b_{n}=na_{n}\mathbb{E}[X]&\text{if \ \ }1<\alpha<2,\\ \beta=0&,a_{n}=\left((c^{+}+c^{-})n\log n\right)^{-1/2}&,b_{n}=na_{n}\mathbb{E}[X]&\text{if \ \ }\alpha=2,\\ \beta=0&,a_{n}=\left(n\sigma^{2}_{X}/2\right)^{-1/2}&,b_{n}=na_{n}\mathbb{E}[X]&\text{if \ \ }\alpha>2.\\ \end{array}$ (30)
@@
$\begin{array}[]{llll}&\frac{(\bar{X}_{N}-\mathbb{E}[X])}{N^{1/\alpha-1}}&\stackrel{{\scriptstyle d}}{{\rightarrow}}\eta_{\alpha}(c^{-}+c^{+})Z&\text{with }Z\sim\mathbf{S}(\alpha,\beta,1,0),\\ &\frac{s^{2}_{N}(X)}{N^{2/\alpha-1}}&\stackrel{{\scriptstyle d}}{{\rightarrow}}\eta_{\frac{\alpha}{2}}(c^{-}+c^{+})Z^{\prime}&\text{with }Z^{\prime}\sim\mathbf{S}(\frac{\alpha}{2},1,1,0),\\ \end{array}$ (31)
@@
Proof. The first result follows directly from Theorem 1 in the case where $21<\alpha<$ 2. To derive the second result, consider the distribution of $Y=X^{2}$. The cumulative distribution function of $Y$ is given by:
@@ Appendix B Correlation distribution in attractive SU(4) Hubbard model
As a direct generalization of the discussion in Section III.1, we can extend the prediction of characteristic exponents to the attractive SU($2N$) Hubbard model. In this case, when approaching the nodal line, the determinant of a single flavor is expected to vanish linearly with the perpendicular distance $x_{\bot}$ from a configuration $\bm{x}$ to the nodal line: $w^{\alpha}(\bm{x})\propto x_{\bot}$, where $\alpha=1,\dots,2N$ denotes the flavor index. In this regime, the Green’s function elements diverge as $G_{ij}\propto 1/x_{\bot}$. The total weight scales as $w(\bm{x})=[w^{\alpha}(\bm{x})]^{2N}\propto x_{\bot}^{2N}$. If an observable is computed as $O_{\text{loc}}=G_{ij}^{m}\propto x_{\bot}^{-m}$ for some integer $m$—for instance, $m=2$ for a pair correlation and $m=4$ for a quartet correlation—then the probability distribution of the local observable behaves as $\mathbb{P}(O_{\text{loc}})\propto|O_{\text{loc}}|^{-1-\frac{2N+1}{m}}$, leading to a characteristic exponent $\alpha=\frac{2N+1}{m}$. Specifically, in the SU(4) case, the charge-$4e$ correlation is associated with a characteristic exponent $\alpha=5/4$, indicating a more severe infinite variance problem than in the SU(2) case. In contrast, the charge-$2e$ correlation has a larger exponent $\alpha=5/2>2$, which places it outside the infinite variance regime.

### v2

@@ III.3 Application to the SU(4) Hubbard model: charge-$4e$ correlations
We next present a real application of the exact bridge link method to a challenging problem, namely the attractive SU(4) Hubbard model. This model is of particular interest for studying quartet condensation, also known as charge-$4e$ superconductivity [52, 53, 54, 55, 56, 57]. Unlike the conventional charge-$2e$ superconducting state, characterized by a pair order parameter involving two creation operators, the charge-$4e$ state is described by the quartet order parameter. Computing the charge-$4e$ correlation function is challenging due to both infinite variance and a large noise-to-signal ratio — particularly because the long-range correlation signal may remain small even in the ordered phase. As we illustrate below, the new method makes it possible to perform “clean” simulations to tackle this problem.
@@ V Acknowledgments
[…]
@@
[…]
@@
Proof. The first result follows directly from Theorem 1 in the case where $1<\alpha<2$. To derive the second result, consider the distribution of $Y=X^{2}$. The cumulative distribution function of $Y$ is given by:
@@ Appendix B Correlation distribution in attractive SU(4) Hubbard model
As a direct generalization of the discussion in Section III.1, we can extend the prediction of characteristic exponents to the attractive SU($2N$) Hubbard model. In this case, when approaching the nodal line, the determinant of a single flavor is expected to vanish linearly with the perpendicular distance $x_{\bot}$ from a configuration $\bm{x}$ to the nodal line: $w^{\alpha}(\bm{x})\propto x_{\bot}$, where $\alpha=1,\dots,2N$ denotes the flavor index. In this regime, the Green’s function elements diverge as $G_{ij}\propto 1/x_{\bot}$. The total weight scales as $w(\bm{x})=[w^{\alpha}(\bm{x})]^{2N}\propto x_{\bot}^{2N}$. If an observable is computed as $O_{\text{loc}}=G_{ij}^{m}\propto x_{\bot}^{-m}$ for some integer mm—for instance, $m=2$ for a pair correlation and $m=4$ for a quartet correlation—then the probability distribution of the local observable behaves as $\mathbb{P}(O_{\text{loc}})\propto|O_{\text{loc}}|^{-1-\frac{2N+1}{m}}$, leading to a characteristic exponent $\alpha=\frac{2N+1}{m}$. Specifically, in the SU(4) case, the charge-$4e$ correlation is associated with a characteristic exponent $\alpha=5/4$, indicating a more severe infinite variance problem than in the SU(2) case. In contrast, the charge-$2e$ correlation has a larger exponent $\alpha=5/2>2$, which places it outside the infinite variance regime.